You are presented with three urns. Each urn has 10 marbles. In addition, each urn holds only black or red marbles. Urn 1 has 3 red marbles. Urn 2 has 8 red marbles. Urn 3 has 5 red marbles. You randomly pick 1 marble from 1 of the three urns. Calculate the probability that it is a black marble. Round answer to the nearest hundredth.


Answer by stanbon(75887)

You are presented with three urns. Each urn has 10 marbles. In addition, each urn holds only black or red marbles. Urn 1 has 3 red marbles. Urn 2 has 8 red marbles. Urn 3 has 5 red marbles. You randomly pick 1 marble from 1 of the three urns. Calculate the probability that it is a black marble. Round answer to the nearest hundredth.
--
total # of marbles: 30
total # of black marbles: 7 + 2 + 5 = 14
---
P(slect a black) = 14/30 = 7/15
===================
cheers,
Stan H.